Circular No. 8401

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2004 S 3, S/2004 S 4, AND R/2004 S 1 C. C. Porco, CICLOPS, Space Science Institute, Boulder; and the

Cassini Imaging Science Team (cf. IAUC 8389) report the discovery of two objects, S/2004 S 3 and S/2004 S 4, orbiting within the F ring of Saturn. S/2004 S 3 orbits just outside the outer strand of the F ring, interior to the orbit of Saturn XVII (Pandora), the outer F-ring shepherd. S/2004 S 4 orbits within the inner F ring strand. S/2004 S 3 was observed in an imaging sequence specifically designed to search for new satellites in the F-ring region and consisting of 19 narrow-angle-camera clear-filter images taken over 2.3 hr on June 21. S/2004 S 4 was observed in a sequence designed to study the photometric behavior of Saturn’s rings taken five hours later and consisting of four narrow-angle-camera images taken 40 s apart in the green, red, blue, and near-infrared broadband filters (Porco et al. 2004, Space Sci. Rev., in press). A low-residual single orbit that crosses the F ring can be fitted to the motions of both objects, but dynamical considerations currently make it unclear whether these objects are one and the same body, and whether they are satellites or clumps. If satellites, the objects are ≈ 4–5 km across. Also, a new ring of diffuse material, R/2004 S 1, has been seen in Cassini images of the F-ring region in the same orbit as the satellite Saturn XV (Atlas).

SUPERNOVAE 2004eg, 2004eh, AND 2004ei Three apparent new supernovae have been discovered on CCD images

taken by J. Young (2004eg; Table Mountain 0.6-m f/16 reflector + R and I filters) and by T. Boles (2004eh and 2004ei; cf. IAUC 8399).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004eg Sept. 1.488 4^{h}28^{m}08.^{s}26 +21^{o}39′18′′.3 19.5 20′′ W, 1′′.4 S
2004eh Sept. 8.003 2 26 11.82 +27 36 24.1 17.1 7′′.8 W, 11′′.6 N
2004ei Sept. 8.074 3 41 32.69 +37 13 30.2 17.9 4′′.3 E, 2′′.9 S

Approximate unfiltered CCD magnitudes from the respective discoverers (plus LOSS/KAIT magnitudes by W. Li, who independently found SN 2004eg): SN 2004eg in UGC 3053, Feb. 5.12, [20.0 (Li); Aug. 20.48, 18.4 (Li); Sept. 1.497, I = 19.0; 2.409, R = 19.5; 6.45, 18.6 (Li); 8.445, R = 19.5. SN 2004eh in UGC 1892, 2003 Oct. 19, [19.5; Nov. 27, [19.5; 2004 Sept. 9.001, 18.5. SN 2004ei in UGC 2817, Jan. 9, [19.5; Feb. 22, [19.5; Sept. 8.998, 17.9. None of the new objects is present on Sky Survey images dating to 1988–1991 (limiting red mag 20.8–21.5; limiting blue mag 21.5–22.5).

2004 September 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green